\begin{thebibliography}{99}
\scriptsize



\bibitem{AH89}
 
 Appel K., Haken W., Every planar map is four colorable,
 A.M.S. Contemp. Math. 98 (1989) 236-240.
 \bibitem{BHKM}
  Becker J., Hewitt J., Kaul H., Maxfield M.,  Mudrock J., Spivey D., Thomason S., 
 Wagstrom T., The DP Color Function of Joins and Vertex-Gluings of Graphs, Discrete Math. 345:11
 (2022), 113093.

\bibitem{Be73} Berge C., Graphs and Hypergraphs, North-Holland, Amsterdam, 1973.

\bibitem{BK19}
 Bernshteyn A., Kostochka A.,
DP-colorings of hypergraphs, European Journal of Combinatorics 78 (2019) 134–146. 

\bibitem{Bi12}
Birkhoff, G.D., A determinant formula for the number of ways of coloring a map. Ann. Math. 14 (1912) 42-
46.
\bibitem{BŁ07}
Borowiecki M., Łazuka E., On chromaticity of hypergraphs, Discrete Mathematics 307 (2007) 1418-1429.

\bibitem{CW78}
  Chao C.Y., Whitehead Jr.E.G., On chromatic equivalence of graphs, Y. Alavi, D.R. Lick, (Eds.), Theory and Applications of Graphs, Lecture
Notes in Mathematics, vol. 642, Springer, Berlin, 1978, pp. 121–131.

\bibitem{DK24}
 Dahlberg S., Kaul H., Mudrock J., An algebraic approach for counting DP-3-colorings of
sparse graphs, European J. Combin. 118 (2024) 103890.





\bibitem{Do93}
 Dohmen K., Chromatische Polynome von Graphen und Hypergraphen, Ph.D. thesis, Düsseldorf, 1993.
\bibitem{DY22}
 Dong F.M, Yang Y., DP color functions versus chromatic polynomials, Adv. Appl. Math. 134
(2022) 102301.




%\bibitem{Go34}

% L. Goeritz, Bemerkungen zur knotentheorie,  Abh. Math. Sem. Univ. Hamburg 10 (1) (1934) 201-210.

\bibitem{DL22}
  Durhuus B., Lucia A., Recursion relations for chromatic coefficients for graphs and
 hypergraphs, Discussiones Mathematicae Graph Theory 42 (2022) 101–121.
\bibitem{DP18}
 %Dvo$\rm\breve{r}\acute{a}$k Z., 
 Dvo\v{r}\'{a}k Z.,
 Postle L., Correspondence coloring and its application to list-coloring planar graphs without cycles of lengths 4 to 8, J. Comb. Theory, Ser. B 129 (2018) 38-54.
 
\bibitem{ERT}
 Erd\H{o}s P., Rubin A.L., Taylor H., Choosability in graphs, Congr. Numer. 26 (1979) 125-127.

\bibitem{HKLMSS} 
 Halberg C., Kaul H., Liu A., Mudrock J.A., Shin P.,  Thomason S., On polynomial representations of the DP color function: theta graphs and their generalizations, J. Comb. 15 (2024) 451-478.
\bibitem{KM20}
 Kaul H., Mudrock J., Combinatorial Nullstellensatz and DP-coloring of Graphs, Discrete
Math. 343 (2020) 112115.


\bibitem{KM21}
 Kaul H., Mudrock J.A., On the chromatic polynomial and counting DP-colorings of graphs, Adv. 
Appl. Math. 123 (2021) 103-121. 
\bibitem{KMSS}

 Kaul H., Mudrock J., Sharma G., Stratton Q., DP-coloring the Cartesian Products of
Graphs, J. Graph Theory 103 (2023) 285-306.

\bibitem{LY24}
 Li Z., Yang Y., Bounds for DP color function and canonical labelings, Graphs Combin. 40 (2024) 65 19pp.
\bibitem{MT21}
 Mudrock J.A., Thomason S., Answers to two questions on the DP color function, Elect.
J. Combin., 28 (2021) 2.24, 14pp.
\bibitem{PS23}
 Postle L., Smith-Roberge E., Exponentially Many Correspondence Colourings of Planar and
Locally Planar Graphs, arXiv:2309.17291v1 (preprint), 2023.
\bibitem{Th09} 
 Thomassen C., The chromatic polynomial and list colorings, J. Comb. Theory, Ser. B 99 (2009)
474–479.
\bibitem{To98}
  Tomescu I., Chromatic coefficients of linear uniform hypergraphs, J. Combin. Theory Ser.
 B 260 (1998) 229–235.
\bibitem{To09}
 Tomescu I., Some properties of chromatic coefficients of linear uniform hypergraphs,
Graphs Combin. 25 (2009) 639–646.
 \bibitem{Tr14}
 Trinks M., A note on a Broken-cycle Theorem for hypergraphs, Discuss. Math. Graph
Theory 34 (2014) 641–646.

\bibitem{Vi76}
 Vizing V.G., Coloring the vertices of a graph in prescribed colors, Diskret. Analiz. no.29, Metody Diskret. Anal. v Teorii Kodovi Skhem 101 (1976) 3-10.
\bibitem{WQY}
  Wang W., Qian J., Yan Z., Colorings versus list colorings of uniform hypergraphs, J.
 Graph Theory 95 (2020) 384–397.

 \bibitem{ZD17}
  Zhang R.X., Dong F.M., Properties of chromatic polynomials of hypergraphs not held
 for chromatic polynomials of graphs, European Journal of Combinatorics, 64 (2017) 138–
 151.

 \bibitem{ZD20} 
 Zhang R.X., Dong F.M., Zero-free intervals of chromatic polynomials of hypergraphs,
 Discrete Math. 343 (2020) 112–134.
 




\end{thebibliography}

\documentclass[a4paper,latexcad,twoside,11pt]{article}
\usepackage{amssymb,amsmath} 
\usepackage{graphicx}
\usepackage{epstopdf}

\usepackage{ntheorem}
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}

\newtheorem{proposition}{Proposition}

\newtheorem{definition}{Definition}

\newenvironment{proof}{{
		\bf Proof\quad}}{\hfill $\hfill\square$\par}

\newtheorem{problem}{Problem}
\def\dfrac{\displaystyle\frac}

\usepackage{color}

\newcommand \red[1]
{\textcolor{red}{#1}}

\usepackage{float} 

\newcommand \equ[2]
{\begin{equation}\label{#1}
		#2
	\end{equation}
}

\newcommand \eqn[2]
{\begin{eqnarray}\label{#1}
		#2
	\end{eqnarray}
}

\newcommand \Bracket[1]
{
	\left	( {#1} \right )
}

\usepackage{stmaryrd}

\newcommand \brk[1]
{
	\llbracket #1\rrbracket
}

\def \hyh {{\mathcal H}}
\def \setf {{\mathcal F}}

\def \dom {\mbox{dom}} 

\def \N {{\mathbb{N}}}

\begin{document}
\title{ DP color functions of hypergraphs }
\author{Ruiyi Cui$^1$, 
Liangxia Wan$^1$\thanks{
		Emails: lxwan@bjtu.edu.cn. },  
Fengming Dong$^2$\thanks{
	Email: fengming.dong@nie.edu.sg}
 \\
 \small $^1$School of Mathematics and Statistics,
Beijing Jiaotong University, Beijing $100044$, China
\\
\small $^2$National Institute of Education, Nanyang Technological University, Singapore
}

\date{}
\maketitle

\abstract{
	In this article, we introduce the DP color function 
	$P_{DP}(\hyh,k)$ 
	of a hypergraph $\hyh$ for $k\in \mathbb N$, based on the DP coloring
	introduced by Bernshteyn and Kostochka, which is the minimum value of $P_{DP}(\mathcal H,\mathcal F)$ 
	all $k$-fold covers $\mathcal F$ of $\mathcal H$. It	
	is an extension 
	of its chromatic polynomial 
	$P(\hyh,k)$ with the property
	that $P_{DP}(\hyh,k)\le P(\hyh,k)$
	for all positive integers $k$.
We obtain an 
	upper bound for $P_{DP}(\hyh,k)$
	when $\hyh$ is a connected 
	$r$-uniform hypergraph for $r\ge 2$, 
	and the upper bound is attained 
	if and only if $\hyh$ is a 
$r$-uniform hypertree. 
	We also show that
$P_{DP}(\hyh,k)= P(\hyh,k)$ holds 
when $\hyh$ is a 
$r$-uniform hypertree or a unicycle linear $r$-uniform hypergraph with odd cycle for $r\ge 3$. These conclusions coincide with  
the known results of graphs.

}
\vskip 2mm
\noindent{\small{\bf Mathematics Subject Classiﬁcation} 05C15, 05C30, 05C65}
\vskip 2mm
\noindent{\bf keywords} Chromatic polynomial$\cdot$ Hypergraph coloring$\cdot$ DP-coloring$\cdot$ DP color function

\section{Introduction}

\noindent

In this paper, each hypergraph is nonempty and simple unless otherwise noted. 
Let $\mathbb{N}$ denote the set of positive integers, and  let 
$\brk{k}$ denote the set $\{1,2,\cdots,k\}$ for $k\in \mathbb{N}$. 
For a hypergraph $\mathcal{H}$,
let $V(\hyh)$ and $E(\hyh)$ denote the 
vertex set and edge set of $\hyh$,
respectively.
For any $k\in \N$, a proper $k$-coloring 
of $\hyh$ is a mapping $f:V(\hyh)\mapsto \brk{k}$
such that for any edge $e\in E(\hyh)$,
there exist a pair of vertices $u,v\in e$
such that $f(u)\ne f(v)$.
This definition clearly includes the
special  case 
that $\hyh$ is a graph. 
The concept of vertex-coloring was
extended to list-coloring 
 independently by Vizing \cite{Vi76} and Erd\H{o}s, Rubin, and Taylor \cite{ERT} in the 1970s. Dvo\v{r}\'{a}k and Postle
~\cite{DP18}
 further  extended list-coloring to DP-coloring (or correspondence coloring) in order to prove that every planar graph without cycles of lengths $4$ to $8$ is $3$-choosable in 2015.

For any graph $G$,  its 
{\it chromatic polynomial}, 
denoted by $P(G,k)$,
is the function counting 
distinct $k$-colorings of $G$. 
This function, indeed a polynomial
in $k$, 
 was introduced by Birkhoff in 1912 with the hope of proving
the Four Color Conjecture \cite{Bi12}. 
This function for a graph is naturally 
extended to that for a hypergraph.
Readers can refer to \cite{DL22, To98, To09, Tr14, WQY, ZD17, ZD20}
for recent works on chromatic 
polynomial.
For any $k\in \N$, 
the list color function 
$P_l(G,k)$ of $G$ is the minimum value 
of $P(G,L)$ over all $k$-assignments
$L$, where  $P(G,L)$ 
is the number of $L$-coloring of $G$.
A survey paper on $P_l(G,k)$ was 
provided 
by Thomassen \cite{Th09}.

\def \seth {{\cal C}}

DP-coloring is a generalization of list coloring,
and a formal definition is given below.
For a graph $G$, a {\it cover} of $G$ is an ordered pair 
$\seth = (L, H)$, 
where $H$ is a graph and 
$L$ is a mapping from $V(G)$ to the power set of $V(H)$
satisfying the four conditions below:
\begin{enumerate}
	\renewcommand \theenumi {(\roman{enumi})}
	
	\item  the set $\{L(u) : u \in V (G)\}$ is a partition of 
$V(H)$;

\item for every $u\in V (G)$, 
$H[L(u)]$ is a complete graph;

\item  if $u$ and $v$ are non-adjacent vertices in $G$,
then $E_H (L(u), L(v))=\emptyset$; and 

\item  for each edge $uv\in E(G)$,  
$E_H (L(u), L(v))$ is a matching.
\end{enumerate}

An {\it $\seth$-coloring} of $G$ is 
defined to be 
an independent set $I$ of $H$ with $|I|=|V(G)|$. 
Clearly,  
$|I\cap L(u)| = 1$ holds for each $u \in V (G)$
and each $\seth$-coloring $I$ of $G$. 
Let $P_{DP}(G,\seth)$ be the number of
$\seth$-colorings of $G$.
A cover $\seth = (L, H)$  of $G$ is called 
an {\it $m$-fold} cover if 
$|L(u)| = m$ for each $u\in V (G)$.
For each $m\in \N$,
let $P_{DP}(G,m)$ be the minimum value 
of $P_{DP}(G,\seth)$ over 
all $m$-fold covers $\seth$ of $G$,
and $P_{DP}(G,m)$ is called 
the {\it DP color function} of $G$,
introduced by 
Kaul and Mudrock~\cite{KM21}. 

In ~\cite{KM21} and  \cite{MT21},
Kaul and Mudrock
showed that for any $g\ge 3$,
there exists 
a graph $G$ with girth $g$ such that $P_{DP}(G, k) < P(G, k)$ holds
for sufficiently large $k$,
and for any graph $G$, 
$P_{DP}(G\vee K_1, k) 
= P(G\vee K_1, k)$ holds for sufficiently large $k$,
where $H_1\vee H_2$
is the graph with vertex set 
$V(H_1)\cup V(H_2)$ and edge set 
$E(H_1)\cup E(H_2)\cup \{xy: x\in V(H_1), y\in V(H_2)\}$.
 For each edge $e$ in $G$, let $l(e) = \infty$ if $e$ is a bridge of $G$, and let $l(e)$ be the length of a shortest cycle in $G$ containing
$e$ otherwise. Dong and Yang \cite{DY22} 
showed that if $l(e)$ is even for some edge $e$ in $G$, then $P_{DP}(G,k) < P(G,k)$ holds for sufficiently large $k$.
They also  proved that $P_{DP}(G,k)=P(G,k)$ 
holds for sufficiently large $k$
if $G$ contains a spanning tree $T$ such that 
for each $e\in E(G)\setminus E(T)$, $l(e)$ is odd and $e$ is contained in a cycle $C$ of length $l(e)$ with the property that $l(e') < l(e)$ for each $e'\in E(C)\setminus (E(T)\cup {e})$. 
Some other recent works 
on DP-color functions 
are referred to \cite{BHKM, DK24, 
	HKLMSS, KMSS, LY24}.

In this article, we introduce the DP color function of a hypergraph.
We obtain an upper bound 
for the DP color function of a 
$r$-uniform hypergraph $\hyh$,
 extending  the 
analogous result of the DP function of a  graph due to Kaul and Mudrock \cite{KM21}. 
We show that this upper bound is attained if and only if  $\cal H$ is a $r$-uniform hypertree for $r\ge 2$.
Corollary $4$ in \cite{KM21} is the 
special case $r=2$ in our result. 
We also find a  formula for 
$P_{DP}(\hyh,k)$ when $\hyh$ is 
a unicycle linear $r$-uniform hypergraph for $r\ge 3$. 
This result extends the Theorem $11$ in \cite{KM21}.

This paper is organized as follows.
In Section $2$, we mainly provide 
some known results 
on the expression of chromatic polynomials of some special uniform hypergraphs, and the definition of 
$k$-fold covers of a hypergraph.   
 In Section 3, we introduce the DP color function of a hypergraph and show several inequalities of the DP color functions of a hypergraph. 
 In Section 4, 
 we prove that the DP color function of any connected $r$-uniform hypergraph $\mathcal H$ equals to the upper bound if and only if $\cal H$ is a $r$-uniform hypertree for $r\ge 2$. 
 Moreover, the explicit formula for the DP color function of any unicycle linear $r$-uniform hypergraph is provided for $r\ge 3$.  Several problems are proposed in Section $5$.

\section{Preliminaries}

In this section, we state some concepts \cite{Be73} and review some related results about the chromatic polynomial of a hypergraph \cite{BK19, BŁ07}.

\subsection{Chromatic polynomials of some hypergraphs}

\vskip 3mm

A {\it hypergraph} $\hyh = (X,E)$ consists of a finite non-empty set
$X$ and a subset $E$ of $2^X$ 
(i.e., the power set of $X$) with 
$|e|\ge 2$ for each $e\in E$.
 A hypergraph $\hyh$ is 
 said to be {\it simple} if 
 $e_1\not\subseteq e_2$ for any two 
 distinct edges $e_1$ and $e_2$ in $\hyh$, 
is {\it $r$-uniform} if $|e| = r$ for each 
$e\in E$, and  is {\it linear} if 
$|e_1\cap e_2|\le 1$ for any two 
distinct edges $e_1$ and $e_2$ in $\hyh$. 
The number of edges in $\hyh$ 
containing a vertex $v$ is its degree $d_{\mathcal H}(v)$.
A {\it cycle} $C$ of length $l$ for $l\ge 3$ in $\mathcal H$ is a subhypergraph consisting of  
$l$ distinct edges
$e_1, e_2,\cdots, e_l$ such that the
vertex $v_i\in e_i\cap e_{i+1}$ exits 
for each $i\in \brk{l-1}$, 
and $v_l\in e_1\cap e_{l}$,
and 
cycle $C$ is
 said to be {\it elementary} if 
 $d_{C}(v_i)=2$ for each 
 $i\in \brk{l}$ and 
 $d_{C}(v)=1$ for any other
 $v\in \cup_{i\in \brk{l}}e_i$.
 
 For any edge $e$ in $\hyh=(X,E)$, 
 $\hyh-e$ denotes the hypergraph $(X,E\setminus\{e\})$.
 A {\it path} $P$ in $\mathcal H$ is a subhypergraph consisting of 
 edges
 $e_1, e_2,\cdots, e_t$ in $\hyh$ such that
 $e_i\cap e_{i+1}\ne \emptyset$ 
 for each $i\in \brk{t-1}$.
 $\hyh$ is said to be {\it connected}
 if for any two vertices $u$ and $v$ in $\hyh$, there exists a path $P$ in $\hyh$ 
 such that both $u$ and $v$ are in $P$.
A {\it hypertree} is a connected linear hypergraph without cycles. 
A {\it unicyclic hypergraph} is a connected hypergraph
containing exactly one cycle.

Dohmen calculated the chromatic polynomials of hypertrees as follows \cite{Do93}.
\begin{lemma}\label{lem:2-1}
	If $\mathcal T_{m}^{r}$ is a $r$-uniform hypertree with $m$ edges, where $r\geq 2$ and $m\geq 0$, then $$P(\mathcal T_{m}^{r},k)=k(k^{r-1}-1)^{m}.$$
\end{lemma}

Borowiecki and Łazuka provided the chromatic polynomials of $r$-uniform unicyclic hypergraphs with $m$ edges as follows \cite{BŁ07}.
\begin{lemma}\label{lem:2-3}
	Let $\mathcal H$ be a linear hypergraph. $\mathcal H$ is a $r$-uniform unicyclic hypergraph with $m + p$ edges and a cycle of length $p$ if and only if
	$$P(\mathcal H,k)=(k^{r-1}-1)^{m+p}+(-1)^{p}(k-1)(k^{r-1}-1)^{m}$$
	where $r\geq 3,m\geq 0$ and $p\geq 3$.
\end{lemma}

\subsection{DP-colorings of a hypergraph}

The concepts of $k$-fold covers and DP-colorings of hypergraphs were first introduced by Bernshteyn 
and Alexandr Kostochka~\cite{BK19}.

Each map could be identified 
with its graph which is the set $\{(x,y)|f(x)=y\}$. Let $\varphi:X\rightharpoonup Y$ indicate 
a partial map from $X$ to $Y$  
defined on a subset of $X$ 
with images in $Y$. 
The set of all partial maps from $X$ to $Y$ is denoted by $[X\rightharpoonup Y]$.

Consider a family 
$\mathcal F\subseteq 
[X\rightharpoonup \brk{k}]$ of partial maps, where $k\in \N$.
The domain of $\mathcal F$, denoted by $\dom(\mathcal F)$,  is 
$$
{\rm dom}(\mathcal F) :=\{{\rm dom}(\varphi):\varphi\in \mathcal F\}.
$$
The set ${\rm dom}(\mathcal F)$ 
can be considered as a hypergraph on $X$. 
Each $\varphi \in {\mathcal F}$ is a subset of the Cartesian product $X\times \brk{k}$, and 
$\mathcal F$ can be regarded as the  hypergraph on $X\times \brk{k}$
whose edge set is the set of 
$\{(x,\varphi(x)): x\in \dom(\varphi)\}$
for all $\varphi\in \setf$.
A function $f:X\rightarrow \brk{k}$ {\it avoids} a partial map $\varphi:X\rightharpoonup \brk{k}$ if $\varphi \nsubseteq f$. Given a family $\mathcal F\subseteq[X\rightharpoonup \brk{k}]$ of partial maps, a {\it $(k,\mathcal F)$-coloring} (or simply an {\it $\mathcal F$-coloring}) is a function $f:X\rightarrow \brk{k}$ that avoids all $\varphi \in \mathcal F$.

\begin{definition}
	\label{k-fold}
	{\rm Let $\hyh$ be a hypergraph 
	and $k\in \N$. 
A $k$-fold cover of $\mathcal H$ 
is a family of partial maps $\mathcal F\subseteq [X\rightharpoonup \brk{k}]$ such that
${\rm dom}(\mathcal F
)\subseteq V(\hyh)$ and
if ${\rm dom}(\varphi)={\rm dom}(\psi)$ for distinct $\varphi,\psi\in \mathcal F$, then $\varphi \cap \psi=\emptyset$.}
\end{definition}

 Let $\mathcal F_e=\{\varphi\in {\cal F}: {\rm dom}(\varphi)=e\}$ and let $\mathcal F\setminus\mathcal F_e=\{\varphi\in {\cal F}: {\rm dom}(\varphi)\neq e\}$.
 If $\setf$ is a $k$-fold cover of $\hyh$,
 then, ${\mathcal F}\setminus\setf_e$ is a $k$-fold cover of ${\cal H}-e$.
If $\mathcal F$ is a $k$-fold cover of a hypergraph $\mathcal H$, then,  
for each edge $e$ in $\hyh$,
 $$
 |\setf_e|=|\{\varphi\in {\cal F}: {\rm dom}(\varphi)=e\}|
 \leqslant k. $$  
 \begin{definition}{\rm
 	A hypergraph $\mathcal H$ on a set $X$ is  $k$-DP-colorable if each $k$-fold cover $\mathcal F$ of $\mathcal H$ admits a $(k,\mathcal F)$-coloring $f:X\rightarrow \brk{k}$.}
\end{definition}

\section{Some properties of the DP color function }

In this section we define a natural cover and the DP color function of a hypergraph. An upper bound is given for the DP color function of any $r$-uniform hypergraph for $r\ge 2$. Several inequalities of the DP color function for any hypergraph $\mathcal H$ are provided by applying a bijection between its $\mathcal F$-colorings and  proper colorings of ${\cal H}-e$ for $e\in E(\cal H)$.
\vskip 2mm

\begin{definition} {\rm
Let $\mathcal H$ be a hypergraph with 
 $E(\hyh)=\{e_j: j\in \brk{m}\}$.
For any $k\in  \mathbb{N}$,  define 
$$
\iota_{\mathcal H}(k)
=\left \{\varphi^{(j)}_i:  i\in \brk{k},  j\in \brk{m}\right \},
$$ 
where $\varphi^{(j)}_i$ is the mapping: $e_j\mapsto \{i\}$.
$\iota_{\mathcal H}(k)$ is 
called a natural $k$-cover of $\hyh$. }
\end{definition}

If $\mathcal I$ is the
set of $\iota_{\mathcal H}(k)$-colorings of $\mathcal H$ and $\mathcal J$ is the set of proper $k$-colorings of $\mathcal H$, then the function
$f: \mathcal J\rightarrow \mathcal I$ given by
$$ f(c) = \{(v,c(v)): v\in V
(\mathcal H)\}$$
is a bijection between the $\iota_{\mathcal H}(k)$-colorings  of $\mathcal H$ and its proper $k$-colorings.

\begin{definition}{\rm
Suppose that $\mathcal F$ is a $k$-fold cover of a hypergraph $\mathcal H$ for $k\in \mathbb{N}$. Let $P_{DP}(\mathcal H,\mathcal F)$ be the number of $\mathcal F$-colorings of $\mathcal H$. The DP color
	function denoted by $P_{DP}(\mathcal H,k)$ is the minimum value of $P_{DP}(\mathcal H,\mathcal F)$
all $k$-fold covers $\mathcal F$ of $\mathcal H$.}
\end{definition} 

It is obvious that
for any hypergraph $\mathcal H$ and $k\in \mathbb{N}$
\begin{equation}
	P_{DP}(\mathcal H,k)\le P(\mathcal H,k).
	\label{eq:3-1}
\end{equation}
If $\mathcal H$ is a hypergraph with components: $\mathcal H_1, \mathcal H_2,\cdots,\mathcal H_t$, then
$$
P_{DP}(\mathcal H,k)=\prod_{i=1}^t P_{DP}(\mathcal H_i,k).$$
So, it is enough to consider the connected hypergraphs 
for the study of 
DP color functions of hypergraphs.

\begin{proposition} Suppose 
	$\mathcal H=(X,E)$ is a $r$-uniform hypergraph with $X=\{x_1,x_2,\cdots,x_n\}$ for $r\ge 2$. Then, for each $k\in \mathbb{N}$, $$P_{DP}(H,k)\leq \dfrac{k^{n}(k^{r-1}-1)^{|E(H)|}}{k^{(r-1)|E(H)|}}.$$
	\end{proposition}
\begin{proof} 
The result holds obviously if $\mathcal{H}$ is edgeless. Suppose that $E(\mathcal{H})
=\{e_1,e_2,\cdots,e_m\}$,
where $m\geq 1$.
Notice that $|[X \rightarrow \brk{k}]|=k^{n}$.
We index the elements 
in $[X \rightarrow \brk{k}]$
as $f_{1},f_{2},\cdots, f_{k^{n}}$.

A $k$-fold cover $\mathcal F$ can be partial randomly formed by choosing a color for each $u\in V(\cal H)$ such that $\varphi^{(j)}_p(u)\ne \varphi^{(j)}_q(u)$ with $p\ne q$ for any $p,q\in \brk{k}$, each $j\in \brk{m}$. Denote the corresponding family partial maps by 
$$
\mathcal{F}
=\{\varphi^{(j)}_i: i\in \brk{k}, 
j\in \brk{m}\}, 
$$ 
where 
$\varphi^{(j)}_i: e_j\rightarrow \brk{k}$.
Let $A_{s}$ be the event that 
$f_{s}$ is a $\mathcal F$-coloring of $\mathcal H$ for $s \in [k^{n}]$.
For any $f\in [X\rightarrow \brk{k}]$
 and $i\in \brk{k}$ and $j\in \brk{m}$, 
 the probability that $f$ avoids $\varphi^{(j)}_i$ in the $k$-fold cover 
 is
$$
1-\Bracket{\dfrac{1}{k}}^{r-1}.
$$
Thus, 
$$
P(A_{s})
=\Bracket{1-\Bracket{\dfrac{1}{k}}^{r-1}}
^{m}.
$$
Let $Y_{s}$ be the random variable that is one if $A_{s}$ occurs and zero otherwise.
Let 
$$Y=\sum\limits_{s=1}^{k^{n}}Y_{s}.$$
Obviously, $Y$ is a random variable and $Y$ is equal to the number of $\mathcal F$-colorings of $\mathcal H$. Because of the linearity of expectation, 
$$E(Y)
=E\Bracket{\sum_{s=1}^{k^{n}}Y_{s}}
=\sum_{s=1}^{k^{n}}
\Bracket{1-\Bracket{\frac{1}{k}}^{r-1}}^{m}
=\frac{(k^{r-1}-1)^{m}\cdot k^{n}}{k^{(r-1)m}}.
$$
The result holds. 
\end{proof}

\vskip 2mm
The following result is obvious for a proper coloring of a hypergraph $\mathcal H$.

\begin{lemma}\label{lem:3-1}
	 Suppose that $\mathcal H$ is a hypergraph and 
	 $e=\{v_i: i\in \brk{r}\}\in E(\hyh)$, 
	 where $r\ge 2$.
	 Assume that $v_i$ can be colored arbitrarily for each $i\in \brk{r}$ in $\cal H$ $-e$. 
	 For any $i_1,i_2,\ldots, i_r$, 
	 where 
	 $i_{j}\in \brk{k}$ for each $j\in \brk{r}$, 
	 let $S_{k}^{(i_1,i_2,\cdots,i_{r})}$ be the set of proper $k$-colorings of $\mathcal H-e$ that color $v_{j}$ with $i_{j}$ for each $j\in \brk{r}$.
Then
\begin{enumerate}
	\item[(1)] 
	$t_1:=|S_{k}^{(i,\cdots,i)}|$ is a constant for all $i\in \brk{k}$, and 
	$$
	kt_1=P(\mathcal H-e,k)
	-P(\mathcal H,k);
	$$
\item[(2)] for any given
$i_1,i_2,\cdots,i_{r}\in \brk{k}$, if 
$i_s\ne i_t$ for a pair of $s,t$ with 
$1\le s<t\le r$,   then 
$t_2:=|S_{k}^{(i_1,i_2,\cdots,i_{r})}|$
is a constant, and 
$$
(k^{r}-k)t_2=k(k^{r-1}-1)t_2=P(\mathcal H,k).
$$
\end{enumerate} 
 \end{lemma}

\begin{lemma}\label{lem:3-2}
	Let $\mathcal F$ be a $k$-fold cover of a hypergraph $\mathcal H$ 
	and $k\geq 2$. 
	Suppose that 
	$e=\{v_i: i\in \brk{r}\}\in E(\mathcal H)$, where $r\ge 2$.
Then, 
$\mathcal F'=\mathcal F-\mathcal F_{e}$ is a $k$-fold cover of $\mathcal H-e$,
where 	
$\mathcal F_{e}=
\{ \varphi\in \mathcal F:
\dom (\varphi) =e\}$.
If there is a 
bijection between $\mathcal F'$-colorings of $\mathcal H-e$ and proper $k$-colorings of $\mathcal H-e$, then	
there exists a $k$-fold cover $\mathcal F^*$ of $\hyh$ such that 
\eqn{eq1}
{
	P_{DP}(\mathcal H,\mathcal F)
	&\geqslant  &
	P_{DP}(\mathcal H,\mathcal F^*) 
	\nonumber \\
	&=&
	\min \left \{
	P(\hyh, k), 
	P(\mathcal H-e,k)	-
	\frac{P(\mathcal H,k)}{k^{r-1}-1}
	\right \}.
}

\end{lemma}
\begin{proof}
	Because there is a 
	bijection between $\mathcal F'$-colorings of $\mathcal H-e$ and its $k$-proper colorings, there are exactly  $P(\mathcal H-e,k)$ $\mathcal F'$-colorings of $\mathcal H-e$. 
	If $v_1, v_2, \cdots, v_r$ all have the same color $i\in \brk{k}$, 
	then there are precisely $(P(\mathcal H-e,k)-P(\mathcal H,k))/k$ $\mathcal F$-colorings of $\mathcal H-e$ which
	contain $(v_{j},{i})$ for all $j\in \brk{r}$ by Lemma \ref{lem:3-1}. 
	Similarly, if each $v_j$ is assigned 
	a color in $\brk{k}$
	and there exists 
	$1\le s<t\le r$ such that 
	$v_s$ and $v_t$ are colored differently, 
	then there are precisely $\dfrac{P(\mathcal H,k)}{k(k^{r-1}-1)}$ $\mathcal H'$-colorings of $\mathcal H-e$ which contain $(v_{j},i_{j})$ for all $j\in \brk{r}$.
	
	Since $|\mathcal F_{e}|\leqslant k$, 
$$P_{DP}(\mathcal H,\mathcal F)\geqslant P(\mathcal H-e,k)-{\rm max}\{P(\mathcal H-e,k)-P(\mathcal  H,k),\frac{P(\mathcal H,k)}{k^{r-1}-1}\}.$$
	A $k$-fold cover $\mathcal F^{*}$ of $\mathcal H$ is constructed as follows.
	If $$P(\mathcal H-e,k)-P(\mathcal H,k)\geqslant \frac{P(\mathcal H,k)}{k^{r-1}-1},
	$$ 
	then, for each $i\in \brk{k}$,
	add $\varphi_i$ with $\varphi(v_j)=i$ for each $j\in \brk{r}$ based on $\mathcal F'$.
	If
	$$P(\mathcal H-e,k)-P(\mathcal H,k)<\frac{P(\mathcal H,k)}{k^{r-1}-1},
	$$ 
	then, $\varphi_{i}:e\rightarrow \brk{k}$ for each $i\in \brk{k}$ such that $\varphi_{i}(v_{1})=i+1$ and $\varphi_{i}(v_{j})=i$ for $2\le j\le r$. 
	It is easy to verify that $\mathcal F^{*}$ has the desired property. Therefore the result holds.
\end{proof}

\begin{proposition} \label{prop1}
Suppose that $\mathcal H$ is a hypergraph and
 $e=\{v_i: i\in \brk{r}\}
 \in E(\mathcal H)$,
 where $r\ge 2$. 
 If 
 $P(\mathcal H-e,k)<\dfrac{k^{r-1}}{k^{r-1}-1}P(\mathcal H,k)$,
then 
$$
P_{DP}(\mathcal H,k)<P(\mathcal H,k).
$$
\end{proposition}
\begin{proof}
Let ${\cal H}'={\mathcal H}-e$.
There
is a bijection between the $ \iota_{\mathcal H'}$-colorings  of $\mathcal H'$ and its proper $k$-colorings.
So, Lemma \ref{lem:3-2} implies that there exist an $k$-fold cover $\mathcal F$ of $\mathcal H$ such that
$$
P_{DP}(\mathcal H, \mathcal F)=\min
\left \{P(\mathcal H, k), P(\mathcal H-e,k)-
\frac{P(\mathcal H,k)}{k^{r-1}-1}\right \}.
$$
Since 
$$
P(\mathcal H-e,k)<\frac{k^{r-1}}{k^{r-1}-1}P(\mathcal H,k)
$$
it follows that 
$$P(\mathcal H-e,k)-\dfrac{P(\mathcal H,k)}{k^{r-1}-1}<P(\mathcal H,k).$$
Then
$$P_{DP}(\mathcal H,k)\leqslant P_{DP}(\mathcal H,\mathcal F)=P(\mathcal H-e,k)-\dfrac{P(\mathcal H,k)}{k^{r-1}-1}<P(\mathcal H,k).$$ 
Thus the result is deduced.
\end{proof}

\section{DP color functions of hypertrees and unicycle hypergraphs}

In this section we show that the DP color function of any connected $r$-uniform hypergraph $\mathcal H$ equals to the upper bound if and only if $\cal H$ is a $r$-uniform hypertree for $r\ge 2$. We provide the explicit formula for the DP color function of any unicycle linear $r$-uniform hypergraph for $r\ge 3$.

	\begin{lemma}\label{lem:3-3}
		Suppose that $\mathcal H$ is a hypergraph with 
	$m$	edges 
		and has no cycles for $m\ge 0$. 
		Suppose that $\mathcal F=\{\varphi_i^{(j)}: i\in \brk{k}, j\in \brk{m}\}$ is a $k$-fold cover
		of $\mathcal H$ for $k\ge 2$. 
		Then there is a 
		bijection between $\mathcal F$ and $\iota_{\mathcal {H}}$.
	\end{lemma}
	\begin{proof}
	We prove the result by induction on $m$. It is obvious for $m = 0,1$.
	Assume that the result holds for 
	any positive integer $m$ $(m\ge 2)$. 
	Now consider the case $m$.  Without loss of generality, suppose that $P=e_1,e_2,\cdots,e_l$ is the longest path.  
		Set 
	$$
	\mathcal H' =\mathcal H - {e_l}, \:
	\mathcal F' =\mathcal F \setminus\mathcal F_{e_l}.
	$$
		Then
	$\mathcal F'$ is a $k$-fold cover of $\mathcal H'$. There is a 
	bijection $f'$ between $\cal F'$ and the natural cover $\iota_{\mathcal H'}$ according to the induction hypothesis. For each $i \in \brk{k}$, let 
	$$
	f(\varphi_i^{(j)}(u))=
	\left\{
	\begin{array}{ll}
		f'(\varphi_i^{(j)}(u)),\qquad 
		&\mbox{ if }j\ne l;\\
		i, &\mbox{ otherwise. }
	\end{array} 
	\right.
	$$
	It is easy to verify that $f$ is a bijection between  its $\mathcal F$ and $\iota_{\mathcal H}$. The conclusion is clear for $m$. Therefore the result holds.
	\end{proof}

\begin{lemma}\label{lem:4-1}
	Suppose that a hypergraph $\mathcal H$ contains one and only one cycle
	$C$, and
	$e=\{v_i: i\in \brk{r}\}$ is an edge in $C$, where $r\ge 2$
	and $v_1$ and $v_2$ are the two vertices in $e$ which also appear in other edges of $C$. 
	Let $S_{k}^{(i_1,i_2,\cdots,i_r)}$ be the set of proper $k$-colorings of $\mathcal H-e$ that color $v_{l}$ by $i_l\in \brk{k}$ for all $l\in \brk{r}$, where $k\ge 2$. Then for any given $i_l\in\brk{k}$
	$$
	t_{1}=\frac{1}{k}(P(\mathcal H-e,k)-P(\mathcal H,k))
	$$
	and
	$$t_{2}=\frac{1}{k^{r-1}(k-1)}(k^{r-2}P(\mathcal H,k)+(1-k^{r-2})P(\mathcal H-e,k)),
	$$ 
	where  $t_1:=|S_{k}^{(i_1,i_2,\cdots,i_r)}|$ 
	when $i_1=i_2$, and $t_2:=|S_{k}^{(i_1,i_2,\cdots,i_r)}|$,
	when $i_1\ne i_2$.
\end{lemma}
\begin{proof}
Because $\mathcal H-e$ doesn't contain any cycle, 
$$
kt_{1}=P(\mathcal H-e,k)-P(\mathcal H,k)
$$
and
$$k(k^{r-2}-1)t_{1}+k^{r-1}(k-1)t_2=P(\mathcal H,k).$$
The conclusion is obtained. 
\end{proof}

\begin{lemma}\label{lem:4-2}
	Suppose that $\mathcal F$ is a 
	$k$-fold cover of a hypergraph $\mathcal H$,
	where $k\in \mathbb{N}$. 
	Suppose that  $\mathcal H$ contains one and only one cycle $C$, and 	
	$e=\{v_i: i\in \brk{r}\}$ is an edge in $C$,
	where $r\ge 2$. 
  If there is a 
	 bijection between the set of $\mathcal F^{'}$-colorings
	 and the set of its proper 
	 $k$-colorings of $\mathcal H-e$,
	  where $\mathcal F^{'}=\mathcal F\setminus\mathcal F_e$, then
	  there exists a $k$-fold cover $\mathcal F^{\ast} $ of $\mathcal H$ such that
	 \eqn{eq3}
	 {
	P_{DP}(\mathcal H, \mathcal F)
	&\geqslant &P_{DP}(\mathcal H, \mathcal F^*)
	\nonumber \\
	&=&\min\left \{ 	P(\mathcal H, k),
\dfrac{(k^{r-1}-1)P(\mathcal H-e,k)-
k^{r-2}P(\mathcal H,k)}
{k^{r-2}(k-1)}
		\right\} .
 }
\end{lemma}
\begin{proof}
	Suppose that 
	$\mathcal H_{1}$ is the connected component of $\mathcal H-e$ where $v_1\in V(\mathcal H_{1})$. Assume that $v_1$ and $v_2$ are the two vertices in $e$ which also appear in other edges of $C$. For any given $i_l\in \brk{k}$, let $S_{k}^{(i_1,i_2,\cdots,i_r)}$ be the set of proper $k$-colorings of $\mathcal H-e$ that color $v_l$ by $i_l$ for each $l\in \brk{r}$. 			
	Let $|S_{k}^{(i_1,i_2,i_3,\cdots,i_r)}|=t_{1}$ when $i_1=i_2$, and 
	and let $|S_{k}^{(i_1,i_2,\cdots,i_r)}|=t_{2}$
	when  $i_1\neq i_2$.

	Note that there is a bijection between its $\mathcal F'$-colorings and $k$-proper colorings of $\mathcal H-e$. If $v_{1}$ and $v_2$ have 
	the same color $i\in \brk{k}$
	in one coloring of $\mathcal H-e$, then 
	there are precisely $t_1$ $\mathcal F'$-colorings of $\mathcal H-e$ which contain $(v_1,i)$ and $(v_2,i)$ for $i\in \brk{k}$; otherwise 
	there are precisely $t_2$ $\mathcal F'$-colorings of $\mathcal H-e$ which contain $(v_1,i)$ and $(v_2,j)$ for $i\neq j,i,j\in \brk{k}$.
	Since $|\mathcal F_{e}|\leqslant k$, it follows that
	$$P_{DP}(\mathcal H, \mathcal F)\geqslant P(\mathcal H-e,k)-{\rm max}\{kt_1,kt_2\}.$$	
	By Lemma \ref{lem:4-1} $$t_{1}=\frac{1}{k}(P(\mathcal H-e,k)-P(\mathcal H,k))$$
	and
	$$t_{2}=\frac{1}{k^{r-1}(k-1)}(k^{r-2}P(\mathcal H,k)+(1-k^{r-2})P(\mathcal H-e,k)).$$

	A $k$-fold cover $\mathcal F^{*}$ of $\mathcal H$ is constructed as follows.
	If $t_1\geqslant t_2,$ 
	then for each $i\in \brk{k}$
	add $\varphi_i$ with $\varphi(v_j)=i$ for each $j\in [n]$ based on $\mathcal F'$.
	If
	$t_1<t_2$, 
	then add $\varphi_{i}:e\rightarrow \brk{k}$ for each $i\in \brk{k}$ such that $\varphi_{i}(v_{1})=i+1$ and $\varphi_{i}(v_{j})=i$ for $2\le j\le r$.
	
	It is easy to verify that $\mathcal F^{*}$ has the desired property. Therefore the result holds.
\end{proof}

\begin{theorem}\label{the:4-3}
	Suppose $\mathcal H$ is a unicycle linear $r$-uniform hypergraph with $m+p$ edges, and its unique cycle is of 
	length $p$ for $p\ge 3$, $r\ge 3$ and $m\ge 0$. 
	Then, 
	\begin{enumerate}
		\item[(1)]  If $p$ is odd, then 
		$$
		P_{DP}(\mathcal H,k)=P(\mathcal H,k).
		$$
		\item[(2)] 
		 If $p$ is even, then 
		$$P_{DP}(\mathcal H,k)=(k^{r-1}-1)^{m+p}+(-1)^{p+1}(k^{r-1}-1)^{m}.$$
	\end{enumerate} 
\end{theorem}

\begin{proof}
Without loss of generality,
 suppose that 
 $E(\mathcal H)=
 \{e_j: j\in\brk{m+p}\}$ and $C$ is the only cycle in $\mathcal H$, and 
 $e=\{v_i: i\in\brk{r}\}$ is an edge in $C$,
 where $v_1$ and $v_2$ are the two
 vertices in $e$ which also appear
 in other edges of $C$. 
  Since adding $\varphi_p^{(j)}$ to $\mathcal F$ can only make the number of $\mathcal F$-colorings of $\mathcal H$ smaller, suppose that 
   $\mathcal F=\{\varphi_i^{(j)}: i\in \brk{k}, j\in \brk{m}\}$ is 
  a $k$-fold cover of $\mathcal H$ with $k\ge 2$. 
   Then $\mathcal H-e_1$ is a hypertree.
	By Lemma \ref{lem:3-3} and Lemma \ref{lem:4-2},
	\begin{equation}	\label{eq:4-3}
		P_{DP}(\mathcal H, \mathcal F)\geqslant	
		P(\mathcal H-e_1,k)-{\rm max}\{A,B\}
	\end{equation}
	where $A=P(\mathcal H-e_1,k)-P(\mathcal H,k)$ and $ B=\dfrac{k^{r-1}P(\mathcal H,k)+(k-k^{r-1})P(\mathcal H-e_1,k)}{k^{r-1}(k-1)}.$
	
	By Lemmas \ref{lem:2-1} and \ref{lem:2-3}, 
	$$A=(k^{r-1}-1)^{m+p-1}+(-1)^{p+1}(k-1)(k^{r-1}-1)^m$$
	and
	$$B=(k^{r-1}-1)^{m+p-1}+(-1)^{p}(k^{r-1}-1)^m.$$
	
	If $p$ is odd, then $A> B$. By applying the inequality \eqref{eq:4-3}, 
	$$P_{DP}(\mathcal H,\mathcal F)\ge P(\mathcal H,k).$$
	By the inequality \eqref{eq:3-1},
	$P_{DP}(\mathcal H,k)= P(\mathcal H,k).$
	If $p$ is even, then $A< B$. By applying the inequality \eqref{eq:4-3}, 
	$$P_{DP}(\mathcal H,\mathcal F)\ge (k^{r-1}-1)^{m+p}+(-1)^{p+1}(k^{r-1}-1)^{m}.$$
	By Lemma \ref{lem:4-2}, there is $\mathcal F^*$ such that 
	$$P_{DP}(\mathcal H,\mathcal F^*)=(k^{r-1}-1)^{m+p}+(-1)^{p+1}(k^{r-1}-1)^{m}.$$
	Therefore, 
	$$P_{DP}(\mathcal H,k)=(k^{r-1}-1)^{m+p}+(-1)^{p+1}(k^{r-1}-1)^{m}.$$
\end{proof}

\begin{lemma} \label{lem:4-4}
	Let	$\mathcal T_m^{r}$ is a $r$-uniform hypertree with $m$ edges for $r\ge 2$. If $m\geqslant 0$, then
	$$P_{DP}(\mathcal T_{m}^{r},k)=P(\mathcal T_{m}^{r},k).$$ 
\end{lemma}
\begin{proof}
	It is obvious for $m=0,1$. Suppose that $\mathcal F$ is any $k$-fold cover of $\mathcal T_m^{r}$ and suppose that $e\in E({\cal T}_m^{r})$. 
	Lemmas \ref{lem:3-2}-\ref{lem:3-3}
	imply that
	
	$$P_{DP}(\mathcal T_{m}^{r},{\cal F})\geqslant P(\mathcal T_{m}^{r}-e,k)-\max\{P(\mathcal T_{m}^{r}-e,k)-P(\mathcal T_{m}^{r},k),\frac{P(\mathcal T_{m}^{r},k)}{k^{r-1}-1}\}.$$
	
	By Lemma \ref{lem:2-1},
	
	$$P(\mathcal T_{m}^{r}-e,k)-P(\mathcal T_{m}^{r},k)=k^{r-1}(k^{r-1}-1)^{m-1}-(k^{r-1}-1)^{m}=(k^{r-1}-1)^{m-1}$$
	and	
	$$\frac{P(\mathcal T_{m}^{r},k)}{k^{r-1}-1}=(k^{r-1}-1)^{m-1}.$$
	
	So $$P_{DP}(\mathcal T_{m}^{r},F)\geqslant P(\mathcal T_{m}^{r},k).$$
	Because $$P_{DP}(\mathcal T_{m}^{r},k)\leqslant P(\mathcal T_{m}^{r},k),$$
	the result is deduced. 
\end{proof}
The case of $r=2$ in the following result implies Corollary $4$ in \cite{KM21}.
\begin{theorem}
	Let	$\mathcal H$ is a connected $r$-uniform hypergraph with $n$ vertices and $m$ edges for $r\ge 2$. 
	Then, for any $k\in \N$,
	\equ{eq4}
	{
	P_{DP}(\mathcal H,k)= \dfrac{k^{n}(k^{r-1}-1)^{m}}
	{k^{(r-1)m}}
}
	if and only if $\mathcal H$ is a hypertree.
\end{theorem}
\begin{proof}
If the hypergraph is a hypertree, then
(\ref{eq4}) holds 
by Lemma \ref{lem:4-4}. Conversely, let suppose that $\mathcal H$ is a connected $r$-uniform hypergraph with $n$ vertices and $m$ edges such that
(\ref{eq4}) holds.

Let $A=\dfrac{k^{n}(k^{r-1}-1)^{m}}{k^{(r-1)m}}$.
If $n=(r-1)m$, then 
$A=(k^{r-1}-1)^{m}$. By 
Theorem \ref{the:4-3}, 
$$
P_{DP}(\mathcal H,k)<(k^{r-1}-1)^{m}.
$$ 
If $n\le (r-1)m-1$, then $A\le \dfrac{1}{k}(k^{r-1}-1)^{m}.$

	Because $\dfrac{1}{k}(k^{r-1}-1)^{m}$ isn't an integer for $k\ge 2$,
$$
P_{DP}(\mathcal H,k)<A.
$$
If $n=(r-1)m+1$, then $\mathcal H$ is a $r$-uniform hypertree.
By the given conditions on $\hyh$,
we have $n\le (r-1)m+1$. 
Therefore the conclusion holds. 
\end{proof}

\section{Further study}

In this section, some problems are proposed whose solutions could make progress on the similar problems in \cite{KM21,Th09}.

\begin{problem}{\rm For which connected hypergraphs $\mathcal H$ does $P_{DP} (\mathcal H, k) = P(\mathcal H, k)$ for every $k \ge 2$?}
\end{problem}
Kaul and Mudrock proved that the answer is affirmative for a tree, a unicycle graph of an odd cycle, chordal and $C_{2p+1}\oplus C_{2q+1}$ \cite{KM20}. 
In the article, we show that the answer is affirmative for a $r$-uniform hypertree and a connected unicycle linear $r$-uniform hypergraph with an odd cycle.

\begin{problem}{\rm
	For a hypergraph $\mathcal H$ with $P_{DP}(\mathcal H, k_0) = P(\mathcal H, k_0)$ for some $k_0\ge \chi(\mathcal H)$, is it true for all $m\ge m_0$ that
	$$P_{DP}(\mathcal H, k) = P(\mathcal H, k)?$$} 
\end{problem}

Kaul and Mudrock provided the affirmative answer of the following problem for a graph \cite{KM20}. 

\begin{problem}
	{\rm For every hypergraph $\mathcal H$, does there exist $p, N\in \mathbb N$ such that $P_{DP} (\mathcal H\vee K_p, k) =P(\mathcal H \vee K_p, k)$ whenever $k\ge N$?}
\end{problem}

\begin{problem}{\rm
Let $\chi_{DP}(\mathcal H)$ denote the smallest positive integer $k$ such that $\cal H$ admits a $\cal F$-coloring for every $k$-fold cover $\cal F$ of a hypergraph $\cal H$. Is it true that if 
$k > \chi(\mathcal H)$,
then
$P_{DP}(\mathcal H,k) > 1$?}
\end{problem}
Thomassen asked if there exists a graph $G$ and 
a $k > 2$ such that $P_l(G, k) = 1$. Clearly, one could make progress on this question by 
showing $P_{DP} (G, m) > 1$ for certain $G$ and $m\in \mathbb N$.

\begin{problem} {\rm For which connected hypergraphs $\mathcal H$ does $P_{DP} (\mathcal H, k) < P(\mathcal H, k)$ for some $k\in \mathbb{N}${\rm ?} Let $P_{DP} (\mathcal H, k) \approx P(\mathcal H, k)$ denote a hypergraph with $P_{DP} (\mathcal H, k) < P(\mathcal H, k)$ and $P_{DP} (\mathcal H, k) = P(\mathcal H, k)$ for enough $k$. Moreover, for which connected hypergraphs $\mathcal H$ does $P_{DP} (\mathcal H, k) \approx P(\mathcal H, k)$?}
\end{problem}
Readers are referred to several results of graphs for the problem above in \cite{DY22, KM21, MT21}.

\begin{problem}
{\rm	For any hypergraph $\mathcal H$ with $n$ vertices, if $P_{DP}(\mathcal H,k)<P(\mathcal H,k)$, is there a constant $\lambda$ such that  for sufficiently large $k$ 
	$$
	P(\mathcal H, k)-P_{DP} (\mathcal H, k) \le\lambda k^{n-2}?
	$$}
\end{problem}

\vskip 5mm
\noindent{\bf Acknowledgements }
\vskip 5mm

This work is supported by National Natural Science Foundation of
China under Grant
No. 12371340.

\end{document}